# Particle Trajectory Representation Learning with Masked Point Modeling

### v2

@@ Abstract
Effective self-supervised learning (SSL) techniques have been key to unlocking large datasets for representation learning. While many promising methods have been developed using online corpora and captioned photographs, their application to scientific domains, where data encodes highly specialized knowledge, remains in its early stages. We present a self-supervised masked modeling framework for 3D particle trajectory analysis in Time Projection Chambers (TPCs). These detectors produce globally sparse ($<1$% occupancy) but locally dense point clouds, capturing meter-scale particle trajectories at millimeter resolution. Starting with PointMAE, this work proposes volumetric tokenization to group sparse ionization points into resolution-agnostic patches, as well as an auxiliary energy infilling task to improve trajectory semantics. This approach – which we call Point-based Liquid Argon Masked Autoencoder (PoLAr-MAE) – achieves 99.4% track and 97.7% shower classification F-scores, matching that of supervised baselines without any labeled data. While the model learns rich particle trajectory representations, it struggles with sub-token phenomena like overlapping or short-lived particle trajectories. To support further research, we release PILArNet-M – the largest open LArTPC dataset (1M+ events, 5.2B labeled points) – to advance SSL in high energy physics (HEP). Project site: https://youngsm.com/polarmae.
@@
Figure 1: Illustration of liquid argon time projection chamber (LArTPC) data. (Left.) High energy particles enter the detector medium, depositing energy along it’s trajectory. These depositions are measured by a complex particle imaging system. (Right.) A significant part of particle trajectory reconstruction is semantic segmentation, whereby each individual occupied voxel is classified as coming from a track-like particle  , electromagnetic shower  , delta ray  , or Michel electron  . A single event can contain both overlapping trajectories and complex interactions in a small amount of space, as illustrated in the two zoomed in portions of the event.
@@ 1 Introduction
Self-supervised learning (SSL) has emerged as a cornerstone of modern computer vision, achieving state-of-the-art results across diverse domains while dramatically reducing reliance on labeled data by using pretrained models as foundation models (Bommasani et al., 2021). By leveraging intrinsic data properties, SSL frameworks learn latent representations that encode rich semantic information, often rivaling fully supervised approaches in downstream tasks. This paradigm shift is especially promising in scientific domains where labeled data is scarce, costly, or impractical to obtain – such as high-energy physics experiments. Indeed, foundation models have steadily made their way into science as an important and active area of research in areas of biology (Jumper et al., 2021; Hou et al., 2024), chemistry (Irwin et al., 2022; Ross et al., 2022), astrophysics (Lanusse et al., 2023; Dillmann et al., 2024), climate science (Lam et al., 2023; Nguyen et al., 2023; Kochkov et al., 2023), mathematics (Trinh et al., 2024; Xin et al., 2024), applied physics (McCabe et al., 2023; Herde et al., 2024), and others.
Time projection chambers (TPCs) (Rubbia, 1977), a modern successor to cloud and bubble chambers, are a crucial detector technology used in a variety of domains to detect and understand the properties of neutrinos (Abi et al., 2020), perform direct dark matter detection (Mount et al., 2017), and understand the structure and properties of atomic nuclei. TPCs operate on the principle of tracking charged particles as they traverse a gas- and/or liquid-filled chamber. As charged particles move through the chamber, they ionize the detector medium, generating a trail of electrons. An electric field within the chamber, often created by a high-voltage cathode, propels these electrons towards an anode where their position and arrival time are precisely recorded. By analyzing this data, the TPC reconstructs the particle’s trajectory and energy deposited in three dimensions, providing valuable insights into its properties and interactions from which it is a result of. An example of TPC data is visualized in Figure 1.
A distinguishing feature of time projection chamber (TPC) data compared to conventional datasets is its extreme sparsity: more than 99% of voxels in a detector volume record no signal in a single event. Particle trajectories manifest as intricate geometric structures – cascading curved 1D manifolds embedded in 3D space. These structures, which vary by TPC type, can span lengths on the order of $\mathcal{O}(1~{}\text{m})$, while modern liquid argon TPCs (LArTPCs) achieve spatial resolutions of $\mathcal{O}(1~{}\text{mm})$. Distinct particle types produce characteristic ionization topologies: tracks, such as those from protons, pions, and muons, appear as long, nearly straight trajectories with minimal curvature. Electromagnetic particles (electrons, positrons, photons) generate shower-like topologies comprising numerous small trajectory fragments. Specific sub-structures include Michel electrons—slightly curved trajectories branching from the Bragg peak of a muon track, resulting from muon decay—and delta rays, electrons produced via interactions between track particles and the liquid argon medium, which appear exclusively along tracks. As illustrated in Fig. 1, these morphologies result in a large diversity of complexity in particle imaging data. Tracks and showers dominate the signal voxels ($>$99.99%), making their accurate discrimination critical for physics analyses. On the other hand, identifying secondary particles like Michel electrons and delta rays can provide insights into the properties of their parent track particles, are thus also important to understand.
Traditional approaches using voxel- and graph-based deep neural networks have achieved significant success in processing data generated from liquid argon TPCs (LArTPCs) (Drielsma et al., 2021), the type of TPC employed in neutrino physics experiments. However, these models rely on labeled datasets derived from physics simulations and suffer from domain shift when applied to real data due to inherent simulation imperfections. This challenge remains a major obstacle to deploying deep learning models in experimental particle physics. Furthermore, the most effective models to date incorporate strong inductive biases rooted in domain science knowledge, and are thus not well-equipt to learn strong representations directly from data. In this paper, we explore adopting PointMAE for analyzing LArTPC image data, aiming to address two key objectives: (1) developing and demonstrating the effectiveness of self-supervised learning (SSL) for our dataset, including masked modeling (Vincent et al., 2010; Pathak et al., 2016; Chen et al., 2020a; Bao et al., 2021; Xie et al., 2022; Hou et al., 2022; He et al., 2021; Pang et al., 2022; Yu et al., 2022), and (2) evaluating whether the SSL approach can compete with or surpass current state-of-the-art models that benefit from inductive bias. We select PointMAE (Pang et al., 2022) as our base model due to its natural ability to handle sparsity while remaining resolution-agnostic, enabling potential applications of these methods to other science datasets.
Inspired by the success of SSL in computer vision and the supervised SPINE (Drielsma et al., 2021) framework for TPC data, this work investigates the potential of self-supervised representation learning for particle trajectory analysis. We aim to bridge the gap between computer vision and particle physics, showcasing how foundational CV methodologies can be adapted to high-energy physics challenges, and how new strategies must be created to deal with the idiosyncratic data modality and precision requirements of particle physics experiments.
Building upon the Point-MAE framework (Pang et al., 2022), we introduce a series of small architectural modifications tailored to our specific data modality, demonstrate that this model – which we name PoLAr-MAE – can effectively learn semantically meaningful representations of particle trajectories, and show that fine-tuning on a 10x less data can yield results comparable to SPINE. Notably, we propose a novel volumetric tokenization approach inspired by naïve voxelization, as well as an auxiliary per-point energy regression task to enhance semantic learning. Using masked modeling as the pretext task, we achieve per-token classification F-scores of 99.4% and 97.7% for identifying tracks and showers. These results are comparable to fully supervised approaches. Additionally, our method attains comparable performance to the state-of-the-art in fine-tuning for dense semantic segmentation of tracks and showers. However, we observe that very fine-grained semantics are not effectively captured with our approach, which poses challenges for downstream dense predictions on particle types that span only one or two tokens (Michel electrons and delta rays ).
Our results demonstrate that self-supervised learning on point-cloud representations can achieve competitive performance in downstream tasks, unlocking the potential of vast, unlabeled raw TPC datasets and paving the way for scalable, cross-experiment analysis of particle trajectories. Along with this paper, we also publish the largest open simulated dataset of LArTPC images, titled PILArNet-M – which contains more than 1,000,000 events and 5.2 billion individually labeled energy depositions – for use in future studies.
In sum, our contributions are as follows:
  • The successful application of the masked modeling framework for learning rich representations of particle trajectories in LArTPCs.
  • A volumetric approach to tokenizing point clouds, which computes a variable number of patch centers per event with empirical guarantees for minimizing missed and duplicated points within patches.
  • An evaluation baseline for pre-training and fine-tuning LArTPC foundation models.
  • A large dataset for use in future research and development by the community.
@@ 2 Related Work
LArTPC data analysis largely relies on the SPINE framework (Drielsma et al., 2021) – a state-of-the-art, end-to-end reconstruction chain for LArTPC data, relying on sparse CNNs for voxel-level feature extraction and GNNs for building higher-level particle structures, and serves as a strong fully-supervised benchmark for our study. SPINE contains a cascade of task-specific networks – either sparse convolution networks or graph neural networks – and was trained end-to-end with over 125,000 individual LArTPC events. It is performant, with dense prediction of tracks and showers achieving 97.7% and 99.5% respectively. A goal of this paper is to determine whether our pre-trained model can enjoy similar performance while fine-tuning on a considerably smaller labeled dataset.
Masked modeling: At the heart of any self-supervised learning framework is the pretext task, i.e., the task under which the model will learn good representations of the data. For our purposes, we focus on two types of pretext tasks: masked modeling (Devlin, 2018) and contrastive learning (Chen et al., 2020b; He et al., 2020; Oord et al., 2018; Tian et al., 2020a, b). The masked modeling task involved masking a part of the training sample, and tasking the model with predicting what it did not see. The success of masked image (He et al., 2021) and language (Devlin, 2018) modeling prompted a number of works to propose masked modeling frameworks for point cloud understanding. Point-MAE (Pang et al., 2022) and Point-BERT (Yu et al., 2022) are perhaps the most popular frameworks, of which many models today use as a baseline for comparison. Point-MAE extended image masked autoencoders (MAEs) (He et al., 2021) to the point-cloud regime by explicitly reconstructing point cloud groupings using the Chamfer distance. Point-BERT extended BERT (Devlin, 2018) pre-training to point clouds. In this work, we start with the Point-MAE architecture, as it is the most conceptually simple and does not require pre-training a point cloud tokenizer as in Point-BERT. In contrast, the contrastive learning task involves maximizing the similarity between augmentations of the same training sample while minimizing the similarity of other samples. Though extremely successful in image tasks, we opt to explore contrastive learning approaches for particle trajectory representation learning in future works, just focusing on whether SSL for LArTPC images is possible in the first place.
SSL for particle physics has recently started gaining momentum, with several papers showing promising results for large-scale pretraining. Golling et al. (2024) proposes a BERT-like (Devlin, 2018) masked particle modeling strategy to learn reusable representations on unordered sets of particles by predicting the representations of masked out particles and comparing them to the output of a frozen pretrained VQ-VAE (Van Den Oord et al., 2017) encoder. Re-Simulation-based Self-Supervised Learning (R3SL) (Harris et al., 2024) utilizes a contrastive-learning approach, where the data augmentations are created by intervening mid-simulation to re-run downstream physics processes with the same initial conditions multiple times over, generating multiple consistent realizations of the same events. This work is different these approaches in our data have not been processed in any way to have coarse particle-based representations, nor do we intervene with strong augmentations as in R3SL. Instead, we work with raw energy depositions, with the only data augmentation being 3D rotations of entire events.
@@
3 Dataset
In this section, we provide an overview of the PILArNet-M dataset. This dataset is similar to the original PIPoLAr-MAE dataset (Adams et al., 2020), with the largest change being an order of magnitude increase in events.
@@ 3 Dataset
For each voxel, the dataset provides the energy deposition in MeV along with additional metadata, including:
  • Fragment ID: A unique integer identifier (per image) representing the smallest unit of a “cluster” of energy depositions.
  • Group ID: A unique integer identifier (within a single event) that aggregates particles with different fragment IDs into larger entities, commonly referred to as “particles” in conventional physics analyses.
  • Interaction ID: A unique integer identifier (within an image) assigned to particles originating from the same physical process or creation event, often referred to as an “interaction” or “event.”
@@
  • Semantic Type: An integer identifier indicating the semantic category of the voxel. Categories include showers (ID: 0), tracks (ID: 1), Michel electrons (ID: 2), delta rays (ID: 3), and spurious low-energy depositions (ID: 4).
This comprehensive dataset enables detailed analyses of particle interactions and energy deposition patterns, supporting both semantic segmentation and clustering tasks. For more information on how the dataset was simulated, see (Adams et al., 2020). Further details on the composition of this dataset can be found in Appendix A.
Data Preprocessing:
We remove low-energy depositions [^1] These spurious spheroid-like deposits provide no meaningful information about the event and can be removed with unsupervised clustering methods like DBSCAN (Ester et al., 1996)., and only train on events with greater than 1024 points. The point cloud is normalized to fit within a unit sphere, and we define $\bar{E}$ as the log-transformed energy of each event, shifted and rescaled to the domain $[-1,1]^{3}$ using the following transformation:
$\bar{E}=2\left(\frac{\log_{10}(E+\epsilon)-\log_{10}(\epsilon)}{\log_{10}(E_{\text{max}}+\epsilon)-\log_{10}(\epsilon)}\right)-1,$
where $E$ is the deposited energy, $\epsilon=0.01$, and $E_{\text{max}}=20$. Here, $\epsilon$ and $E_{\text{max}}$ represent the smallest and largest energy values observed in the dataset, respectively.
We normalize events uniformly across the dataset to ensure that voxel values are processed consistently in a physically meaningful manner. Only basic 3D rotations are applied as training augmentations, for other transformations (e.g., scaling, jittering) would result in non-physical events.
Input Features:
The input to our point cloud architecture consists of all non-zero voxels in a single LArTPC image, represented as a list of 3D coordinates with a single energy channel. Each image is parameterized as the set $\mathcal{X}=\{(\textbf{x}_{i},\bar{E}_{i})\}_{i=1}^{N}$, where $N$ is the number of points in an event, $i$ denotes the $i$-th point, $\textbf{x}_{i}$ is the 3D coordinate of the $i$-th point, and $\bar{E}_{i}$ is the log-normalized energy as defined above.
Training and validation sets:
The training set consists of 1,033,377 events containing at least 1024 points. A smaller dataset of approximately 30 events (15,000 tokens) from a cache of 10,471 held-out events are used for validation metrics for pre-training, and another 30 events from the same cache are used for testing. For fine-tuning, we use a variable number of events from the training set, and use all 10,471 held-out events in the validation set for evaluation.
Evaluation metrics:
To evaluate the performance of pre-training on this dataset, we follow previous works to train linear Support Vector Machines (SVMs) to classify individual point groups as related to different semantic IDs. As tokens may contain a number of points containing any number of depositions from different particles, we fit one classifier per class and perform multi-class classification with a One-vs-the-rest (OvR) strategy, i.e., our classifier classifies tokens as containing specific semantic types or not. We report F-scores ($F_{1}$), which are defined as the harmonic mean of the precision and recall of the classifier, i.e. $F_{1}=\frac{2}{\text{recall}^{-1}+\text{precision}^{-1}}$. To evaluate the performance of the dense (point-wise) semantic segmentation fine-tuning, we report F-scores across classes, as well as overall precision for use in comparison to the fully supervised approach.

### v3

@@ Abstract
Effective self-supervised learning (SSL) techniques have been key to unlocking large datasets for representation learning. While many promising methods have been developed using online corpora and captioned photographs, their application to scientific domains, where data encodes highly specialized knowledge, remains a challenge. Liquid Argon Time Projection Chambers (LArTPCs) provide high-resolution 3D imaging for fundamental physics, but analysis of their sparse, complex point cloud data often relies on supervised methods trained on large simulations, introducing potential biases. We introduce the Point-based Liquid Argon Masked Autoencoder (PoLAr-MAE), applying masked point modeling to unlabeled LArTPC images using domain-specific volumetric tokenization and energy prediction. We show this SSL approach learns physically meaningful trajectory representations directly from data. This yields remarkable data efficiency: fine-tuning on just 100 labeled events achieves track/shower semantic segmentation performance comparable to the state-of-the-art supervised baseline trained on $>$100,000 events. Furthermore, internal attention maps exhibit emergent instance segmentation of particle trajectories. While challenges remain, particularly for fine-grained features, we make concrete SSL’s potential for building a foundation model for LArTPC image analysis capable of serving as a common base for all data reconstruction tasks. To facilitate further progress, we release PILArNet-M, a large dataset of 1M LArTPC events. Project site: https://youngsm.com/polarmae.
@@
Figure 1: Illustration of liquid argon time projection chamber (LArTPC) data. (Left.) High energy particles enter the detector medium, depositing energy along their trajectories. These depositions are measured by a complex particle imaging system. (Right.) A significant part of particle trajectory reconstruction is semantic segmentation, whereby each individual occupied voxel is classified as coming from a track-like particle  , electromagnetic shower  , delta ray  , or Michel electron  . A single event can contain both overlapping trajectories and complex interactions in a small amount of space, as illustrated in the two zoomed in portions of the event.
@@ 1 Introduction
Liquid Argon Time Projection Chambers (LArTPCs) are a cornerstone technology in modern experimental neutrino physics, enabling detailed studies of neutrino oscillations, interactions, and searches for physics beyond the Standard Model (Abi et al., 2020). As high-resolution imaging detectors (Rubbia, 1977), LArTPCs capture the three-dimensional trajectories of charged particles produced in neutrino interactions within large volumes of liquid argon. When charged particles traverse the argon, they leave trails of ionization electrons. These electrons drift under a strong electric field and are collected on finely segmented anode planes, allowing for precise reconstruction of particle paths and energy depositions, often achieving millimeter resolution over meter-scale distances (visualized in Fig. 1). The resulting data, while globally sparse ($>99\%$ empty voxels), contains intricate ionization patterns corresponding to different particle types. These topologies are critical for event interpretation.
Minimum ionizing particles like muons, protons, and charged pions produce long, relatively straight tracks characterized by dense, linear ionization trails. In contrast, electrons, positrons, and photons initiate electromagnetic showers, which appear as diffuse, often conically expanding structures composed of many short, scattered track segments resulting from pair production and bremsstrahlung. Beyond these primary topologies, two important secondary structures provide additional physics information. Michel electrons arise from the decay of a stopping muon ($\mu\rightarrow e^{-}~{}\nu~{}\bar{\nu}$) and manifest as short (typically few cm), sometimes slightly curved electron tracks originating specifically from the Bragg peak of a muon track. Delta rays are energetic electrons knocked out of Ar atoms by a high-energy charged particle traversing the medium; they appear as short electron tracks branching off directly from the side of a primary particle track. Accurately reconstructing and identifying these distinct track, shower, Michel, and delta ray topologies from the complex LArTPC images is a critical step for achieving the physics goals of current and future neutrino experiments. An example event containing these complex topologies is shown in the right side of Fig. 1.
The current state-of-the-art approach for LArTPC data reconstruction leverages deep neural networks trained with supervised learning on large, detailed Monte Carlo simulated datasets, as demonstrated by frameworks such as SPINE (Drielsma et al., 2021b). This “simulate $\rightarrow$ supervised train $\rightarrow$ calibrate” paradigm has proven highly effective and is central to ongoing analyses. However, like any approach, it involves certain trade-offs and presents aspects that motivate the exploration of complementary techniques. For instance, developing and validating these supervised models requires significant investment in simulation generation and careful calibration to mitigate potential discrepancies between simulation and real detector data (domain shift) (Acciarri et al., 2017; Adams et al., 2019). Models trained for one detector geometry or condition may not be directly transferable, and managing distinct networks for the multitude of reconstruction tasks (segmentation, clustering, particle ID, etc.) can be complex. The computational resources required for large-scale simulation also remain a consideration for next-generation experiments. These factors, alongside the desire for methods that might offer greater adaptability or uncover unexpected features in data, encourage investigation into alternative machine learning strategies.
Self-supervised learning (SSL) offers a promising path forward. By learning representations from the inherent structure of unlabeled data itself, SSL frameworks, often developed as foundation models (Bommasani et al., 2021), have shown remarkable success in computer vision and other scientific fields (Jumper et al., 2021; Hou et al., 2024; Irwin et al., 2022; Ross et al., 2022; Lanusse et al., 2023; Dillmann et al., 2024; Lam et al., 2023; Nguyen et al., 2023; Kochkov et al., 2023; Trinh et al., 2024; Xin et al., 2024; McCabe et al., 2023; Herde et al., 2024) for their robustness, transferability of representations, and out-of-distribution generalization. In this “pre-train $\rightarrow$ fine-tune” paradigm, a large model is initially pre-trained on a vast corpora of unlabeled data using unsupervised or self-supervised learning to learn efficient representations of the data. Then, the model is fine-tuned in a supervised manner to utilize these learned representations to very quickly adapt to any downstream task with a small amount of data. This trivial adaption to very few examples is often described as ‘few-shot learning’ within ML literature.
SSL’s application to the unique data characteristics and analysis workflows of LArTPCs remains largely unexplored. This paper presents the first study applying masked modeling directly to raw, unlabeled 3D LArTPC point cloud data. We investigate whether SSL can learn meaningful physical representations that enable highly data-efficient downstream performance.
To this end, we develop the Point-based Liquid Argon Masked Autoencoder (PoLAr-MAE). Building upon the Point-MAE framework (Pang et al., 2022) designed for generic point clouds, PoLAr-MAE incorporates modifications tailored to LArTPC data: a novel volumetric tokenization method to effectively group sparse ionization points into meaningful patches, and an auxiliary energy prediction task to focus the model on calorimetric information. The model is pretrained on a large dataset of unlabeled simulated LArTPC events using a masked autoencoding objective, i.e., learning to reconstruct masked portions of particle trajectories.
Our results demonstrate that PoLAr-MAE learns powerful representations that give rise to efficient transfer to pixel-level tasks. Fine-tuning PoLAr-MAE for semantic segmentation achieves performance on par with the fully supervised Sparse UResNet (Graham et al., 2018; Drielsma et al., 2021b) baseline for identifying dominant track and shower topologies, while requiring drastically less labeled data – matching the baseline trained on O(100,000) events using only O(100) labeled events for fine-tuning. This highlights the potential of SSL to dramatically reduce the need for large labeled datasets, a significant advantage in the context of complex detector simulations and potential domain shifts – the so-called ‘sim2real‘ gap. Qualitative analysis reveals that specific attention heads in the transformer encoder learn to focus on individual particle instances, effectively performing emergent instance segmentation without explicit supervision. However, we also find that resolving fine-grained, sub-token phenomena like Michel electrons and delta rays remains challenging for the current architecture, indicating clear avenues for future work.
This work establishes the viability and data efficiency of SSL masked modeling for LArTPC data analysis, paving the way for more robust, scalable, and adaptable reconstruction workflows in current and future experiments like DUNE. To facilitate further research and development by the community in this emerging area, we release PILArNet-M, a large simulated LArTPC dataset, containing over one million events and 5.2 billion labeled energy depositions.
In sum, our main contributions are:
  • The first successful application and adaptation of self-supervised masked modeling for representation learning directly on sparse 3D point cloud data from LArTPCs.
  • Demonstration that SSL-pretrained representations capture physically meaningful structures, evidenced by emergent instance segmentation in attention maps and extreme data efficiency for track/shower classification.
  • Introduction and validation of C-NMS, a volumetric tokenization strategy tailored for sparse particle trajectory data, and demonstration of the benefit of an auxiliary energy prediction task.
  • The public release of PILArNet-M, a large LArTPC simulation dataset, to serve as a benchmark and resource for future ML research in particle physics.
@@ 2 Related Work
LArTPC Data Analysis and Reconstruction. The reconstruction of particle interactions in LArTPCs is a complex task tackled by sophisticated software frameworks. The current state-of-the-art often involves algorithms based on deep neural networks trained in a supervised manner on large Monte Carlo simulations. A prominent example is the SPINE framework (Drielsma et al., 2021b), which employs a combination of sparse convolutional networks (SCNs) (Dominé et al., 2020, 2021) and graph neural networks (GNNs) (Drielsma et al., 2021a) in a multi-task cascade to perform dense segmentation, clustering, and particle identification. SPINE’s Sparse UResNet achieves high performance in semantic segmentation, particularly for track and shower classification (reporting 97.7% and 99.5% precision, respectively, when trained on over 125,000 simulated events), and serves as the primary supervised benchmark for evaluating the downstream task performance of our self-supervised model. A key motivation for our work is to investigate whether SSL can result in representations that are naturally separable and require very little fine-tuning to achieve comparable performance to the fully supervised approach.
Self-Supervised Learning for Point Clouds. Our work leverages recent advances in self-supervised learning adapted for 3D point cloud data. Many successful approaches build upon the Transformer architecture (Vaswani, 2017), originally developed for natural language processing and later adapted for computer vision (Vision Transformer, ViT) (Dosovitskiy, 2020) and point clouds (Guo et al., 2021; Zhao et al., 2021). A powerful SSL paradigm is masked modeling, inspired by its success in language (Bidirectional Encoder Representations from Transformers, BERT) (Devlin, 2018) and vision (Masked Auto-encoder, MAE) (He et al., 2021). This involves masking portions of the input data and training the model to predict the missing content. For point clouds, Point-MAE (Pang et al., 2022) and Point-BERT (Yu et al., 2022) demonstrated the effectiveness of this approach, typically by dividing the point cloud into local patches, masking a majority of them, and using a Transformer encoder-decoder to reconstruct the masked information. We base our PoLAr-MAE model on the Point-MAE architecture due to its conceptual simplicity and strong performance, adapting its patchification and reconstruction targets for the specific nature of LArTPC data, as detailed in Sec. 4. While contrastive learning (Chen et al., 2020; He et al., 2020; Wilkinson et al., 2025) and self-distillation (Oquab et al., 2024; Zeid et al., 2023) methods represent additional major SSL directions, we focus here on masked modeling as a first exploration of SSL for raw LArTPC images.
Representation Learning in High Energy Physics. Self-supervised learning is gaining traction within HEP, although applications often differ significantly based on the physics context (e.g., collider vs. neutrino physics) and the data representation used. For example, (Golling et al., 2024) proposed masked modeling on sets of reconstructed particles (like jets), predicting properties of masked particles using features learned by a VQ-VAE. R3SL (Harris et al., 2024) employed contrastive learning where augmentations were generated by re-simulating physics processes, requiring access to simulation internals and ground truth. Other efforts focus on contrastive learning for particle tracking at colliders, often operating on graph representations of detector hits or using supervised objectives (Lieret & DeZoort, 2024; Miao et al., 2024). Specific to LArTPCs, recent work explored contrastive learning following SimCLR (Chen et al., 2020) aiming for representations robust to detector variations (Wilkinson et al., 2025). Interestingly, that study found that using realistic detector variations (like noise levels or wire efficiencies) as the sole source of augmentation provided insufficient signal for learning powerful representations, indicating that stronger augmentations or pretext tasks are likely necessary for LArTPC data. Geometric Deep Learning (GDL), particularly methods incorporating physical symmetries like SE(3)-equivariance (Satorras et al., 2021; Fuchs et al., 2020), represents another important direction for analyzing 3D physics data, although these architectures can be more complex to implement and train. Our work distinguishes itself by applying unsupervised masked modeling – a strong pretext task involving fine-grained reconstruction – directly to the raw, high-resolution, sparse 3D energy depositions (point clouds) from LArTPCs, without relying on reconstructed physics objects, simulation interventions, or explicit geometric GDL architectures, aiming to learn fundamental representations of particle trajectories in this specific data modality.
@@
3 Dataset and Evaluation
In this section, we provide an overview of the simulated LArTPC dataset used for training and evaluation, PILArNet-M, and outline our evaluation strategy. This dataset represents an expansion of the earlier PILArNet dataset (Adams et al., 2020) through an order-of-magnitude increase in the number of simulated events.
@@ 3 Dataset and Evaluation
For each active voxel, the dataset provides the deposited energy (in MeV) along with metadata labels derived from the simulation truth, enabling detailed analysis:
  • Fragment ID: A unique identifier (per event) for the smallest contiguous clusters of energy depositions, useful for low-level clustering tasks.
  • Group ID: A unique identifier (per event) aggregating fragments belonging to the same physical particle trajectory (e.g., a single muon track), often corresponding to “particles” in physics analyses.
  • Interaction ID: A unique identifier (per event) grouping particles originating from the same primary physics interaction vertex.
@@
  • Semantic Type: An integer classifying the particle type responsible for the energy deposition. Categories include electromagnetic showers (ID: 0), minimum ionizing particle tracks (ID: 1), Michel electrons (ID: 2), delta rays (ID: 3), and low-energy depositions often considered background (ID: 4).
This comprehensive labeling supports evaluating various reconstruction tasks, including the semantic segmentation focused on in this work, as well as potential future studies on particle clustering or interaction vertexing. Further details on the dataset composition and statistics can be found in Appendix A.
3.1 Data Preprocessing
Before input to the network, several preprocessing steps are applied. First, voxels labeled as low-energy depositions (ID: 4) are removed, as these typically represent noise or diffuse background hits not associated with primary particle trajectories  [^1] These spurious deposits provide minimal physics information for the tasks considered here and can often be removed effectively using standard clustering techniques like DBSCAN (Ester et al., 1996). and considerably increase the memory usage of our network. We only consider events containing at least 1024 remaining points after this cut. The 3D coordinates ($\textbf{x}_{i}$) of the remaining points in each event are normalized to lie within a unit sphere centered at the origin. The energy depositions ($E_{i}$) are log-transformed to handle their large dynamic range and then scaled to the interval $[-1,1]$:
$\bar{E}_{i}=2\left(\frac{\log_{10}(E_{i}+\epsilon)-\log_{10}(\epsilon)}{\log_{10}(E_{\text{max}}+\epsilon)-\log_{10}(\epsilon)}\right)-1,$ (1)
where $\epsilon=0.01$ MeV and $E_{\text{max}}=20$ MeV are chosen based on the observed energy range in the dataset. This normalization ensures consistent input scaling for the network. During training, the only data augmentation applied is random 3D rotation of the entire event point cloud; other geometric augmentations like scaling or geometric jittering are avoided as they may correspond to non-physical transformations of the particle trajectories and energy depositions.
3.2 Training and Validation Sets
The full dataset contains 1,033,377 events meeting the preprocessing criteria. This set is used for self-supervised pre-training. A held-out set of 10,471 events is reserved for validation and testing. For evaluating the quality of the pre-trained representations (linear probing), a small random subset of  30 validation events is used. For evaluating the downstream semantic segmentation task after fine-tuning, we train on subsets of the main training set (with varying sizes, down to 100 events) and evaluate performance on the full 10,471-event held-out validation set.
3.3 Evaluation Metrics
We employ two stages of evaluation: one during pre-training, and one during fine-tuning. To assess the semantic information captured by the SSL representations without full fine-tuning, we follow standard practice (Pang et al., 2022; He et al., 2021) and train linear Support Vector Machines (SVMs) on the frozen representations (tokens, i.e., patches of points) produced by the pre-trained encoder. Since each token represents a group of points that may contain multiple semantic types, we train separate One-vs-Rest (OvR) classifiers to predict the presence of each semantic type within a token. Performance is measured using the $F_{1}$-score (harmonic mean of precision and recall: $F_{1}=2/(\text{recall}^{-1}+\text{precision}^{-1})$) for each class. Additionally, after fine-tuning a segmentation head on top of the pre-trained encoder (using varying amounts of labeled data), we evaluate point-wise semantic segmentation performance. We report standard classification metrics, including per-class $F_{1}$-scores and precision, allowing direct comparison with the fully supervised Sparse UResNet baseline. Results are presented in Sec.5, with further detailed results in Appendix E.